OpenStudy (bowtie_man123):

Which is a solution to the equation 5x + 2y = –1? A. (2, –4) B. (1, –3) C. (–2, 5) D. (–1, 4)

OpenStudy (alexandervonhumboldt2):

substitue (x, y) points and check.

OpenStudy (anonymous):

B Is the answer

OpenStudy (bowtie_man123):

Does eveyone agree

OpenStudy (alexandervonhumboldt2):

yes but please no direct answers

OpenStudy (alexandervonhumboldt2):

5*1-2*3=-1 5-6=-1 -1=-1